Freeze and Hot Tap Case History #22
Well Information: While functioning gate valve the stem had been twisted off with the valve half closed. Well was secured with blind rams on the 70mPa BOP stack… Surface casing pressure was 42mPa.
Procedure: Hole volume of fluid was pumped down the casing to lower working pressure to 10 mPa. Gel was then injected by the freeze crew into the surface and production casings. Gel was then frozen using dry ice and freeze was held on for 24 hours. Freeze plug was then pressure tested to 84 mPa, tested OK. Pressure was then bleed off above freeze plug and monitored. Atmospheric test was good. A crane was then used to lift 70 mPa stack and malfunctioning orbit valve off of well. New orbit valve was then installed along with 70 mPa stack. New orbit valve and BOP stack was then pressure tested. Thawing operations were performed by rig crew.
Time on location: 24 hours